OpenStudy (anonymous):

if logb(3^b) = b/2, then b = a. 1/9 b. 1/3 c. 1/2 d. 3 e.9 please show how to get to the solution!

OpenStudy (anonymous):

Use opposite operations until you get to expressions you can evaluate: GEt rid oof the log by taking to exponent of b \[\log_{b} (3^b)=b/2\] \[3^b=b ^{(b/2)}\]

OpenStudy (anonymous):

Can you take it from there?